CS 582 — Database Management System II — Spring 2007

Quiz #7

NAME: We would like to develop a XML document representing information about the university. For that, we need a type that allows for the storing of the following information:

• The university has several departments.

• Each departement is identified by a name, which is a string.

• The university has several employees

• Each employee works in a department, has a unique university ID, and has a name, which is a string. Create a type for this purpose.